The length is 4 units less than 3 times the width. The perimeter is 22 units more than twice the width. What is the length, width, and perimeter?
Answer by macston(5194)

W=width; L=3W-4; P=perimeter=2W+22
P=2(L+W)
2W+22=2(3W-4+W) Divide each side by 2.
W+11=4W-4 Add4 to each side.
W+15=4W Subtract W from each side.
15=3W Divide eachside by 3
5=W ANSWER 1: The width is 5
L=3W-4=3(5)-4=15-4=11 ANSWER 2: The length is 11.
P=2(L+W)=2(11+5)=2(16)=32 ANSWER 3: The perimeter is 32
CHECK:
Perimeter=2(L+W)
2W+22=2(11+5)
2(5)+22=2(16)
10+22=32
32=32